Hudson Co. reports the contribution margin income statement for 2015.
HUDSON CO.
Contribution Margin Income Statement
For Year Ended December 31, 2015
  Sales (10,500 units at $276 each) $ 2,898,000
  Variable costs (10,500 units at $207 each) 2,173,500
  Contribution margin $ 724,500
  Fixed costs 400,200
  Pretax income $ 324,300
2.

If Hudson achieves its target pretax income for 2016, what is its margin of safety (in percent)? (Round your answer to 1 decimal place.)
Answer #1

Contribution margin ratio=Contribution margin/Sales

=724500/2,898,000

=0.25

Breakeven=Fixed expenses/Contribution margin ratio

=400200/0.25

=$1600800

Margin of safety=Total sales-Breakeven sales

=2,898,000-1600800

=1297200

Margin of safety%=Margin of safety/Total sales

=1297200/2,898,000

=44.8%(Approx).